The ratio of Sam's age to Hank's age is 5 to 3. If the sum of their ages is 24, how old is Hank?
Answer by ewatrrr(24785)

 
Hi,
Let x represent the constant of proportionality
5x + 3x = 24
8x = 24
x = 3
Hank is 5*3 = 15
CHECKING our answer*** Sam is 9
15 + 9 = 24
Wish You the Best in your Studies.